OpenStudy (anonymous):

for questions 5 and 6, add or subtract 5. (m2-m-4)+(m-5) 6. (4x^2-2x-1)-(-3x^3+2)

OpenStudy (er.mohd.amir):

Q 5. m2-m-4+m-5 =m2-9 Q 6. 4x2-2x-1+3x3-2=3x3+4x2-2x-3

OpenStudy (anonymous):

thank you so much!